# Mechanism and Emergence of Stacked Attention Heads in Multi-Layer Transformers

### v2

@@ 5 Theoretical Analysis of Information Flow
In this section, I provide a formal proof that the retrieval problem requires a minimum number of transformer layers depending on the number of retrieval steps $D$. I model the information flow between different positions during self-attention under the following simplifying assumptions:
@@
Definition 1. Let’s consider all relevant input positions ordered by their reachability from the last token following exactly the paths in Figure 3 (e.g., ${\bm{x}}_{0}{\bm{x}}_{0}{\bm{x}}_{1}{\bm{x}}_{1}{\bm{x}}_{2}{\bm{x}}_{2}\ldots$ ). Let’s denote this sequence of positions as ${\bm{s}}_{i}$ for $i\in\{0,1,\ldots,2D\}$. In other words, ${\bm{s}}_{2k}$ and ${\bm{s}}_{2k+1}$ are the positions of the second and first occurrences of ${\bm{x}}_{k}$, respectively. For example, ${\bm{s}}_{0}$ is the last position in the input sequence, ${\bm{s}}_{1}$ is the first occurrence of ${\bm{x}}_{0}$, ${\bm{s}}_{2}$ is the second occurrence of ${\bm{x}}_{1}$, ${\bm{s}}_{2D}$ is the position of the target token ${\bm{x}}_{D}$, and so on.
[…]
Theorem 1 shows that the retrieval problem with $D$ steps cannot be solved by transformers with less than $\log_{3}(2D)$ layers. This result is a lower bound that does not take into account the limitations of network capacity, causal masking, or training dynamics. In practice, I expect the number of required layers to be even higher.
@@ 7 Implicit Curriculum & Number of Layers
My initial experiments suggest that the implicit curriculum (IC) plays a crucial role in the successful learning of the retrieval problem. To better quantify this effect, I conduct two comprehensive sets of experiments, one for each formulation (IC and non-IC). For each formulation, I train 64 transformers with 1 to 8 layers (8 transformers for each number of layers). We plot the final validation loss averaged across all runs with the same number of layers in Figure 4 (left).
@@ 8 Reverse-Engineering the Circuits Learned
To understand the mechanism learned by transformers to solve the retrieval problem, I train three transformers (denoted as A, B, and C) with 12 layers and only one attention head per layer on the retrieval problem with $D=3$. I then manually reverse-engineer the circuits learned by the transformers by studying their attention maps. The uncovered circuits are depicted in Figure 5. I describe my reverse-engineering process in detail in Appendix D.
@@ 8.2 Results
I find that transformers A, B, and C achieve a validation mean squared error of less than $0.01$. By studying the attention maps in the trained transformers, I observe that most attention heads do not perform any useful computation. Only a few attention heads are responsible for the information flow. Their behavior is easily interpretable (see Appendix B).
I manually reverse-engineer the entire circuits learned by the three transformers, which are depicted in Figure 5. I perform extensive validations of the circuits using ablations. My reverse-engineering process is described in detail in Appendix D.
@@ 9 Emergence of Attention Heads during Training
To better understand how the retrieval heads emerge during training, I train a 24-layer transformer (denoted as Transformer D) on the retrieval problem with $D=4$ and IC. I manually reverse-engineer the circuits learned by Transformer D following the same procedure described in Appendix D. Afterward, I measure the attention during training for each attention path in the reverse-engineered circuit.
@@ 9.2 Results
For every checkpoint, I measure the average attention for each attention path in the reverse-engineered circuit. I approximate the attention between checkpoints using linear interpolation. I display the plots for each attention path in Appendix C. Finally, I show the first epoch when the average attention goes above $0.5$ for each attention path (Figure 7, right).
@@
    Finally, two more heads emerge on layers 23 and 24 that reuse the circuit formed by heads 14, 21, and 22 to retrieve the tokens D and E, respectively. This drives down the last two partial losses.
@@ 11 Related Work
By studying two-layer transformers, Olsson et al. (2022) uncover a mechanism termed induction head that, given an input sequece ab…a, can predict b. One possible use of an induction head is sequence copying, but it can also perform more high-level functions such as translation. The induction head is formed by two stacked attention heads. The first head copies into the residual stream of b the value of the previous token a. The second head is then able to attend to the token b and copy it into the residual stream of the final token. Reddy (2023)  explains the emergence of the induction head during training by the sequential learning of three nested logits enabled by an implicit curriculum.
@@ 12 Conclusion
In this work, I introduced the retrieval problem, a simple task that requires transformers to retrieve information from multiple positions in the input sequence. I showed that the retrieval problem requires a certain number of layers to be solved. By training transformers on a minimal formulation of the task, I found that transformers solve the task using a mechanism that resembles an induction head. I found that this mechanism emerges gradually with the help of an implicit curriculum, starting with an induction head and then adding more heads one by one.
@@ Appendix A Retrieval and Conditional Retrieval Prompts
Below I provide one complete example for each of the retrieval and conditional retrieval formulations used in the paper. The examples are generated using the same programs used for benchmarking the large language models. Each example consists of a prompt, a correct answer, and acceptable answers. Acceptable answers are used to filter out incoherent answers by repeateadly sampling from the model until an acceptable answer is found.
@@ A.3 Kingdoms formulation
Silvanians believes in celestianism.
Novarians believes in harmonianism.
Aurorans believes in elysianism.
Florinians believes in luminism.
@@ A.5 Relatives formulation (conditional retrieval)
Appendix B Attention Maps
@@ B.2 Transformer D
Appendix C Attention Emergence in Transformer D
@@
Appendix D Procedure for Reverse-Engineering Circuits
As can be seen in Appendix B, the attention maps for each attention head in the transformers contain clear patterns that can be be manually identified without the need for any additional tools. For this reason, I decided to manually reverse-engineer the circuits, while validating them thoroughly using ablations to ensure their correctness.

### v4

@@ 5 Theoretical Analysis of Information Flow
In this section, I theoretically establish that solving the retrieval problem requires a minimum number of transformer layers that grows logarithmically with the number of retrieval steps $D$. I model the information flow between different positions during self-attention under the following simplifying assumptions:
@@
Based on this simplified model of the attention mechanism, we can prove the following result:
Theorem 1. The last position in the sequence cannot retrieve the embedding vector ${\bm{x}}_{D}$ of the target token with $t$ transformer layers if $t<\log_{3}(2D)$.
This implies that at least $\log_{3}(2D)$ transformer layers are required to solve a retrieval problem with $D$ retrieval steps. Note that this result is a lower bound that does not take into account the limitations of network capacity, causal masking, or training dynamics. In practice, we can expect the number of required layers to be even higher.
I defer the complete proof to Appendix B. The intuition behind this result comes from the fact that the last position in the sequence cannot retrieve the target token ${\bm{x}}_{D}$ without retrieving everything in-between (i.e., ${\bm{x}}_{1}$, ${\bm{x}}_{2}$, $\ldots$, ${\bm{x}}_{D-1}$). However, it is possible to show that the number of retrieved tokens can grow at most by a factor of 3 after each attention layer, hence the logarithmic growth of the number of required layers.
@@ 7 Implicit Curriculum & Number of Layers
My initial experiments suggest that the implicit curriculum (IC) plays a crucial role in the successful learning of the retrieval problem. To better quantify this effect, I conduct two comprehensive sets of experiments, one for each formulation (IC and non-IC). For each formulation, I train 64 transformers with 1 to 8 layers (8 transformers for each number of layers). I plot the final validation loss averaged across all runs with the same number of layers in Figure 4 (left).
@@ 8 Reverse-Engineering the Circuits Learned
To understand the mechanism learned by transformers to solve the retrieval problem, I train three transformers (denoted as A, B, and C) with 12 layers and only one attention head per layer on the retrieval problem with $D=3$. I then manually reverse-engineer the circuits learned by the transformers by studying their attention maps. The uncovered circuits are depicted in Figure 5. I describe my reverse-engineering process in detail in Appendix E.
@@ 8.2 Results
I find that transformers A, B, and C achieve a validation mean squared error of less than $0.01$. By studying the attention maps in the trained transformers, I observe that most attention heads do not perform any useful computation. Only a few attention heads are responsible for the information flow. Their behavior is easily interpretable (see Appendix C).
I manually reverse-engineer the entire circuits learned by the three transformers, which are depicted in Figure 5. I perform extensive validations of the circuits using ablations. My reverse-engineering process is described in detail in Appendix E.
@@ 9 Emergence of Attention Heads during Training
To better understand how the retrieval heads emerge during training, I train a 24-layer transformer (denoted as Transformer D) on the retrieval problem with $D=4$ and IC. I manually reverse-engineer the circuits learned by Transformer D following the same procedure described in Appendix E. Afterward, I measure the attention during training for each attention path in the reverse-engineered circuit.
@@ 9.2 Results
For every checkpoint, I measure the average attention for each attention path in the reverse-engineered circuit. I approximate the attention between checkpoints using linear interpolation. I display the plots for each attention path in Appendix D. Finally, I show the first epoch when the average attention goes above $0.5$ for each attention path (Figure 7, right).
@@
    Finally, two more heads emerge on layers 23 and 24 that reuse heads 14, 21, and 22 to retrieve tokens D and E, respectively. This drives down the last two partial losses.
@@ 11 Related Work
By studying two-layer transformers, Olsson et al. (2022) uncover a mechanism termed induction head that, given an input sequence ab…a, can predict b. One possible use of an induction head is sequence copying, but the authors argue that it can also perform more high-level functions such as translation. The induction head is formed by two stacked attention heads. The first head copies into the residual stream of b the value of the previous token a. The second head is then able to attend to the token b and copy it into the residual stream of the final token. Reddy (2023) explains the emergence of the induction head during training by the sequential learning of three nested logits enabled by an implicit curriculum.
@@ 12 Conclusion
In this work, I introduced and studied the retrieval problem, a simple task that requires transformers to retrieve information from multiple positions in the input sequence. I showed that the retrieval problem requires a certain number of layers to be solved. By training transformers on a minimal formulation of the task, I found that transformers solve the task using a mechanism that resembles an induction head. I found that this mechanism emerges gradually with the help of an implicit curriculum, starting with an induction head and then adding more heads one by one.
@@ Appendix A Retrieval and Conditional Retrieval Prompts
Below I provide one complete example for each of the retrieval and conditional retrieval formulations used in the paper. The examples are generated using the same programs used for benchmarking the large language models. Each example consists of a prompt, a correct answer, and acceptable answers. Acceptable answers are used to filter out incoherent answers by repeatedly sampling from the model until an acceptable answer is found.
@@ A.3 Kingdoms formulation
Silvanians believe in celestianism.
Novarians believe in harmonianism.
Aurorans believe in elysianism.
Florinians believe in luminism.
@@ A.5 Relatives formulation (conditional retrieval)
Appendix B Proof of Theorem 1 (minimum number of layers)
In this section, I provide the complete formal proof for Theorem 1 stated in Section 5. Recall the two assumptions underlying our simplified model of the attention mechanism:
Definition 1. Let’s consider all relevant input positions ordered by their reachability from the last token following the paths of shared information before any attention layers, exactly as depicted in Figure 8 (e.g., ${\bm{x}}_{0}{\bm{x}}_{0}{\bm{x}}_{1}{\bm{x}}_{1}{\bm{x}}_{2}{\bm{x}}_{2}\ldots$ ). Let’s denote this sequence of positions as ${\bm{s}}_{i}$ for $i\in\{0,1,\ldots,2D\}$. In other words, ${\bm{s}}_{2k}$ and ${\bm{s}}_{2k+1}$ are the positions of the second and first occurrences of ${\bm{x}}_{k}$, respectively. For example, ${\bm{s}}_{0}$ is the last position in the input sequence, ${\bm{s}}_{1}$ is the first occurrence of ${\bm{x}}_{0}$, ${\bm{s}}_{2}$ is the second occurrence of ${\bm{x}}_{1}$, ${\bm{s}}_{2D}$ is the position of the target token ${\bm{x}}_{D}$, and so on.
[…]
Appendix C Attention Maps
@@ C.2 Transformer D
Appendix D Attention Emergence in Transformer D
@@
Appendix E Procedure for Reverse-Engineering Circuits
As can be seen in Appendix C, the attention maps for each attention head in the transformers contain clear patterns that can be be manually identified without the need for any additional tools. For this reason, I decided to manually reverse-engineer the circuits, while validating them thoroughly using ablations to ensure their correctness.